# VIRES: Video Instance Repainting with Sketch and Text Guidance

### v2

@@ Abstract
footnotetext: Equal contribution; Corresponding author.
@@ 7.1 Variations of Sequential ControlNet
Ours (VIRES) Variation_1 Variation_2 Variation_3 block_1 Conv (3, 72) Conv (3, 36) Conv (3, 72) Conv (72, 72) Conv (36, 36) Conv (72,72) Down (72, 144) Down (3,64) Down (36, 72) Down (72, 144) block_2 Conv (144, 144) Conv (64, 64) Conv (144,144) Conv (144, 288) Conv (64, 128) Conv (144, 288) Conv (288, 288) Block (288, 288) Block (128, 128) Block (72, 72) Block (288, 288) Down (288, 576) Down (128, 128) Down (72, 288) Down (288, 576) block_3 Block (576, 576) Block (128, 256) Block (288, 288) Block (576, 576) Down (576, 1152) Down (256, 256) Down (288, 1152) Down (576, 1152) block_4 Block (1152, 1152) Conv (1152, 1152) Block (256, 256) Conv (1152, 1152) Conv (1152, 1152) Conv (1152, 1152) Conv (1152, 1152) Conv (1152, 1152) Conv (1152, 1152) Conv (256, 1152) Conv (1152, 1152) Conv (1152, 1152)

### v3

# VIRES: Video Instance Repainting via Sketch and Text Guided Generation

@@ Abstract
footnotetext: Equal contribution. Now at Beijing Academy of Artificial Intelligence. Corresponding author.
@@ 7.1 Variations of Sequential ControlNet
1 2 3 4 block_1 Conv (3, 72) Conv (3, 36) Conv (3, 72) Conv (72, 72) Conv (36, 36) Conv (72,72) Down (72, 144) Down (3,64) Down (36, 72) Down (72, 144) block_2 Conv (144, 144) Conv (64, 64) Conv (144,144) Conv (144, 288) Conv (64, 128) Conv (144, 288) Conv (288, 288) Block (288, 288) Block (128, 128) Block (72, 72) Block (288, 288) Down (288, 576) Down (128, 128) Down (72, 288) Down (288, 576) block_3 Block (576, 576) Block (128, 256) Block (288, 288) Block (576, 576) Down (576, 1152) Down (256, 256) Down (288, 1152) Down (576, 1152) block_4 Block (1152, 1152) Conv (1152, 1152) Block (256, 256) Conv (1152, 1152) Conv (1152, 1152) Conv (1152, 1152) Conv (1152, 1152) Conv (1152, 1152) Conv (1152, 1152) Conv (256, 1152) Conv (1152, 1152) Conv (1152, 1152)